OpenStudy (anonymous):

Use a graphing calculator to solve the equation -6cos(pi/3)q=4 in the interval from 0 to 2p. Round to the nearest hundredth.

OpenStudy (anonymous):

can you help me @ZeHanz

OpenStudy (zehanz):

Do you have a ti84/ti83?

OpenStudy (anonymous):

No, I've got a ti89 though

OpenStudy (zehanz):

Should work the same way, I guess.. BTW, what is the letter "q" in your equation?

OpenStudy (anonymous):

theta

OpenStudy (zehanz):

Oh, ok! YOu have to go to the screen where you can type Y=.... Then type: \(Y_1=-6\cos(\pi/3*X)\) (the "\(Y_1=\)"-part is already there). Also: \(Y_2=4\) Then go to the WINDOW menu, to set the window variables: \(Xmin=0\\Xmax=2\pi\\Xscl=\pi/6\\Ymin=-6\\Ymax=6\\Yscl=1\\Xres=1\) Now draw the graph (GRAPH-button) You will get a inverted cosine graph and a horizontal line at height 4.

OpenStudy (zehanz):

If you now go to the CALC menu, choose 5: intersect, then 3 times Enter. The intersection found is one of the solutions. Do the same again, but only hit Enter twice (that is to choose the graphs), and then move your cursor to the other intersection and then hit Enter. The second solution will be found.

OpenStudy (zehanz):

Sorry, have to go now, be back later...

OpenStudy (zehanz):

Here's how my screen looked: